OpenStudy (anonymous):

how long will it take using simple interest for $1,200 to amount to $1,650 at 8%, rounding to the nearest tenth.

OpenStudy (anonymous):

1200* 8 * t /100 = 450 => t = 450/96 = 4.7 years...